# Detecting Neurocognitive Disorders through Analyses of Topic Evolution and Cross-Modal Consistency in Visual-Stimulated Narratives

### v2

@@ Abstract
Early detection of neurocognitive disorders (NCDs) is crucial for timely intervention and disease management. Given that language impairments manifest early in NCD progression, visual-stimulated narrative (VSN)-based analysis offers a promising avenue for NCD detection. Current VSN-based NCD detection methods primarily focus on linguistic microstructures (e.g., pauses, lexical diversity), which are potentially linked to bottom-up (stimulus-driven) cognitive processing. While these features illuminate basic language abilities, the higher-order linguistic macrostructures (e.g., thematic or logical development), which may reflect top-down (concept-driven) cognitive abilities, remain underexplored. These patterns are crucial for NCD detection yet challenging to quantify due to their abstract and complex nature. To bridge this gap, we propose two novel dynamic macrostructural approaches: (1) Dynamic Topic Model (DTM) to track topic evolution over time, and (2) Text-Image Temporal Alignment Network (TITAN) to measure cross-modal consistency between speech and visual stimuli. Experimental results validated the efficiency of proposed approaches in NCD detection, with TITAN achieving superior performance both on the CU-MARVEL-RABBIT corpus (F1 = 0.7238) and the ADReSS corpus (F1 = 0.8889). The feature contribution analysis revealed that macrostructural features (e.g., topic variability, topic change rate, and topic consistency) constituted the most significant contributors in the model’s decision pathways, outperforming investigated microstructural features. These findings underscore the critical role of macrostructural patterns in understanding cognitive impairment mechanisms in NCDs.
Index Terms:
Neurocognitive disorder detection, Alzheimer’s Disease, macrostructure, temporal narrative analysis, sequential image storytelling, natural language processing, image processing, multi-modal recognition.
publicationid: pubid: $0$–$00$ YEAR IEEE
@@ I Introduction
Neurocognitive disorders (NCDs)[1, 2], such as Alzheimer’s Disease, are highly prevalent and pose significant challenges to individuals and healthcare systems in our aging society. Early detection of these conditions is critical, as it can allow for timely intervention, potentially slowing disease progression and improving the quality of life for patients[3]. Traditional clinical approaches such as neuroimaging [4, 5], retinal photographs [6], electroencephalography (EEG) [7], cerebrospinal fluid (CSF) analysis [8], and in-person neuropsychological assessments [9] have been widely investigated. However, these approaches can be resource-intensive, invasive, and costly [10, 11]. Consequently, there is a growing emphasis on developing non-invasive, cost-effective, and scalable methods for early NCD detection.
@@
One promising approach is the automatic analysis of speech and language, as communicative impairments are often among the earliest symptoms of cognitive decline [12]. These impairments may include temporal changes in spontaneous speech [13], word finding and retrieval difficulties [14], verbal dysfluencies [15], reduction in structural and thematic organization abilities [16, 17]. Such changes in speech patterns and language use provide valuable insights into underlying cognitive deficits and serve as potential biomarkers for early NCD detection. In light of this, various spoken cognitive tests, such as counting [18] and naming [19], have been designed to discover spoken language correlates for NCD detection. Among these, the visual-stimulated narrative (VSN) task [20, 21], such as describing a single image or telling a story based on image sequences, stands out by employing visual stimuli to elicit rich and spontaneous speech within a naturalistic and holistic framework. Specifically, the VSN task engages processes like viewing, describing, sequencing, and integrating elements, requiring multiple cognitive functions, such as attention, visual comprehension, language fluency, executive planning, memory, and social cognition [22, 23, 24, 25, 26, 27, 28]. Therefore, it offers superior potential for comprehensively evaluating cognitive abilities for NCD detection [29, 12].
@@
Approaches to the analysis of VSNs can be categorized into two complementary aspects  [30]. First, the mainstream approach is microstructural analysis, which examines detailed, localized elements of a narrative, such as linguistic (e.g., lexicon, semantics, syntax) [31, 32] and acoustic (e.g., prosody, phonology) [33, 34] units. This level focuses on how things are produced and primarily reflects bottom-up (exogenous, stimulus-driven) cognitive processing. In contrast, macrostructural analysis evaluates the overarching organization of the narrative units, such as informativeness (overall information accuracy and completeness), coherence (logical flow of events or concepts) and text-visual consistency (alignment with visual stimuli [^1] This includes both depiction per image and connection across images, bridging informativeness and coherence.[35]. This level focuses on what is said and how things are combined, generally representing top-down (endogenous, concept-driven) cognitive integration.
@@
Clinical research indicates that individuals with NCDs often exhibit greater impairment in macrostructuring than microstructuring. While microstructural deficits typically reveal specific language abilities (perception and production), macrostructural deficits more directly reflect impairments in higher-order cognitive abilities (attention, memory, organization, executive function), manifesting as disrupted topic maintenance, incoherent sequencing, and inconsistency to visual stimuli [30]. Despite this, macrostructural analysis remains underexplored, particularly within the context of NCD narratives.
@@
This research gap primarily arises from the inherent difficulty in quantifying macrostructural indicators due to their abstract and conceptual nature. Traditional approaches to quantify coherence rely on manual assessments, which are subjective and labor-intensive [36]. Recent advances in natural language processing (NLP) have introduced automated methods, such as calculating cosine similarities between consecutive textual embeddings or using large language models (LLMs) as evaluators [37, 38]. However, these methods often fall short for NCD narratives. Embedding-based similarities can misinterpret frequent repetitions or fillers as better cohesion and ignore dynamic evolutions of narrative topics over time. LLM-generated scores lack interpretability due to their untraceable decision pathways, undermining reliability in clinical applications.
@@
To quantify the consistency between narratives and visual stimuli, recent studies have used text-visual pretrained models to identify the highest relevance scores between image regions and sentences [39] or pattern-matching with fixed references [40, 41]. However, these methods have limitations: local text-image relevance fails to capture global consistency, and pattern-matching is restricted to fixed-reference tasks, making it unsuitable for creative tasks like storytelling. Therefore, tailored and automated approaches are needed for macrostructural analysis of NCD narratives.
@@
This work proposed to investigate specific cognitive and linguistic challenges by analyzing topical shifts, their temporal dynamics, and the coherence of narratives over time. We presented two novel approaches to macrostructural narrative analysis: Dynamic Topic Models (DTM) for temporal analysis of topical evolution; and Text-Image Temporal Alignment Model (TITAN) for cross-modal (i.e., stimulus images and spoken narratives) alignment analysis. This approach aims to reveal cognitive deficits by identifying narrative impairments associated with NCDs, and to provide a deeper understanding of their impact on communication and cognition. The investigations are conducted using the Rabbit Story task (a VSN task) from the CU-MARVEL corpus (a Cantonese corpus), as well as the single-image description task in the ADReSS corpus (an English corpus). Our experiments validated the effectiveness of dynamic topic consistency as a macrostructure metric (F1=0.65, AUC=0.78). The proposed TITAN model achieved the highest performance (F1=0.72, AUC=0.81) in classification tasks, surpassing those of established microstructure and handcrafted macrostructure feature sets. Moreover, experiments of cross-comparison and regression tasks further demonstrated the effectiveness of the proposed dynamic macrostructure modeling approaches for NCD detection.
@@
The paper is organized as follows: Section II reviews related works, Section III introduces corpora used in this study, Section IV introduces the proposed approaches, Section V presents our experimental settings and results, Section VI analyzes and discusses experimental results, and Section VII concludes this work and prospects future works.
@@ II Related Work
Extensive research has been conducted for NCD detection based on spoken language, leveraging various methodologies to identify biomarkers and assess cognitive impairments. This section reviews pertinent literature encompassing both microstructural and macrostructural narrative analysis, collectively laying the groundwork for enhanced narrative analyses. The term “microstructure” refers to fine-grained elements of narrative production, including lexical, syntax, semantics, phonetics, acoustics, and other smaller linguistic units. In contrast, the term “macrostructure” refers to higher-level aspects of narrative construction, such as coherence, cohesion, thematic development, and the overall organization of narratives. Both microstructural and macrostructural analysis are crucial for a comprehensive understanding of cognitive deficits reflected in NCD narratives. However, existing studies often face limitations in capturing the full spectrum of narrative features, particularly the dynamic aspects of narrative construction. By identifying these limitations, this review highlights the gaps the proposed methods aim to fill, specifically by incorporating temporal dynamics and macrostructural consistency, thus advancing the narrative analysis of NCDs.
@@ II-A Microstructural Analysis in NCD Detection
Individuals with NCDs exhibit speech deficits that impact spoken language across both acoustic and linguistic dimensions. Prior research has explored the various aspects of acoustic and linguistic impairments in NCD speech, encompassing prosody (e.g., increased dysfluencies and longer, more frequent pauses), phonetics (e.g., abnormal Mel Frequency Cepstral Coefficients patterns), phonology (e.g., higher rates of mispronunciation of words), lexicon (e.g., reduced type-token-ratio and elevated stopword ratio), syntax (e.g., reduced parse-tree depth), and semantics (e.g., abnormal patterns in word and sentence embeddings) [31, 42, 33, 43, 44, 13, 45, 46, 32, 47, 48, 40, 49, 50, 51]. These microstructural features are essential for understanding the language abilities of individuals with NCDs, providing the foundation for analysis in this domain.
Recent efforts are devoted to identifying highly correlated and significant biomarkers of NCDs from the aforementioned features [52, 53, 54, 55, 56]. For example, Cho et al. [56] linked higher CSF p-tau [^2] CSF p-tau: Cerebrospinal fluid phosphorylated tau, a NCD-related biomarker indicating the burden of neurofibrillary tangle pathology. levels (signifying greater burden of NCD pathology) to specific acoustic and linguistic features: lower preposition and noun ratios, higher pause rates, and shorter speech segment durations. Such microstructural analysis provides valuable insights into local features, but more comprehensive information may be offered by macrostructural approaches that examine broader narrative structure and contexts. Hence, this work extends prior research by examining VSN’s macrostructure such as topic evolution and cross-modal consistency.
(a) The Cookie Theft picture
(b) The Rabbit Story cartoon
Figure 1: Visual stimuli used for NCD detection: (a) Cookie Theft (depicting two kids stealing cookies while a distracted mother is washing dishes, amidst a complex background) and (b) Rabbit Story (showing a girl and her dog searching for a rabbit and encountering various animals and adventures).
@@ II-B Macrostructural Analysis in NCD Detection
While microstructural analysis highlights specific language production impairments associated with NCDs, macrostructural analysis is equally important as it reveals how cognitive deficits affect higher-order cognitive processes like attention, working memory, planning, and organization. This macrostructural perspective is essential for understanding how NCDs affect the construction of coherent, cohesive, and thematically consistent narratives [57, 58, 59, 30].
Macrostructural analysis in the literature generally focuses on features about cohesion (the use of linguistic devices for structural unity) and coherence (logical flow and overall consistency). Individuals with NCDs tend to produce text with wrong co-referencing, inconsistencies in maintaining the main topic, uncompleted or irrelevant information, and disrupted sequencing of events [60, 61, 62, 63, 64].
Traditional macrostructural features were usually extracted manually based on expert experience, such as readability [65], relationship between a proposition and previous one or the global topic [66], or connection between events, macro-propositions and main information units [67].
Recent advances leverage machine learning and large language models (LLMs) for automatic analyze macrostructure of VSNs. Some studies have quantified global semantic coherence by using cosine similarity of consecutive embeddings [37, 45], while others have assessed pragmatic coherence via BLEU and METEOR scores using references [41]. More recent work has adopted LLM as evaluator to automatically predict macro-descriptors [38].
However, as mentioned in Section I, these methods are unsuitable for NCD spoken narratives: similarity scores are confounded by frequent repetitions, reference-based metrics (BLEU, METEOR) are inapplicable due to non-standardized references, and LLM-based evaluation fails to meet interpretability requirements. Additionally, there has been limited investigation into the temporal dynamics of topical evolution and the cross-modality alignment between visual stimuli and speech in VSN tasks. These gaps highlight the need for more advanced approaches to evaluate NCD narratives from both macrostructural and temporal perspectives, addressing the interplay between evolving topics and the integration of visual and spoken elements.
@@ III Corpora
This study is conducted based on the ADReSS (Alzheimer’s Dementia Recognition through Spontaneous Speech) Corpus [68], as well as the CU-MARVEL-RABBIT (Cognitive Assessment Using Machine LeArning EmpoweRed VoicE AnaLysis – RABBIT Story Part) Corpus. These corpora were selected to enable a comprehensive understanding of VSN patterns for NCDs across different language and narrative types. ADReSS collects English recordings from the Cookie Theft picture description task [69], while CU-MARVEL collects Cantonese narratives from the Rabbit Storytelling task [70].
III-A ADReSS
@@
The ADReSS Corpus, developed by the University of Pittsburgh, contains transcripts of spontaneous speech samples from elderly English-speaking adults. The corpus is designed to facilitate research on age-related cognitive decline, focusing on NCDs (specifically, Alzheimer’s disease and mild cognitive impairment). The neurocognitive assessment tasks in this corpus involve the Cookie-Theft Picture Description, as shown in Fig. 1(a), for both Healthy Control (HC) and NCD groups. This task assesses participants’ ability to recognize the elements and produce a coherent narrative based on their spatial relationships and implied events. This corpus consists of audio recordings and transcriptions of interviews with 156 participants in the Cookie-Theft Picture Description task, balanced partitioned into training (54 HCs and 54 NCDs) and testing (24 HCs and 24 NCDs) subsets.
III-B CU-MARVEL-RABBIT

### v3

@@ Abstract
Early detection of neurocognitive disorders (NCDs) is crucial for timely intervention and disease management. Given that language impairments manifest early in NCD progression, visual-stimulated narrative (VSN)-based analysis offers a promising avenue for NCD detection. Current VSN-based NCD detection methods primarily focus on linguistic microstructures (e.g., lexical diversity) that are closely tied to bottom-up, stimulus-driven cognitive processes. While these features illuminate basic language abilities, the higher-order linguistic macrostructures (e.g., topic development) that may reflect top-down, concept-driven cognitive abilities remain underexplored. These macrostructural patterns are crucial for NCD detection, yet challenging to quantify due to their abstract and complex nature. To bridge this gap, we propose two novel macrostructural approaches: (1) a Dynamic Topic Model (DTM) to track topic evolution over time, and (2) a Text-Image Temporal Alignment Network (TITAN) to measure cross-modal consistency between narrative and visual stimuli. Experimental results show the effectiveness of the proposed approaches in NCD detection, with TITAN achieving superior performance across three corpora: ADReSS (F1=0.8889), ADReSSo (F1=0.8504), and CU-MARVEL-RABBIT (F1=0.7238). Feature contribution analysis reveals that macrostructural features (e.g., topic variability, topic change rate, and topic consistency) constitute the most significant contributors to the model’s decision pathways, outperforming the investigated microstructural features. These findings underscore the value of macrostructural analysis for understanding linguistic-cognitive interactions associated with NCDs.
@@ I Introduction
Neurocognitive disorders (NCDs)[1, 2], such as Alzheimer’s Disease, are highly prevalent and pose significant challenges to individuals and healthcare systems in our aging society. Early detection of these conditions is critical, as it facilitates timely intervention and disease management [3]. Traditional clinical approaches such as neuroimaging [4, 5], retinal photographs [6], electroencephalography (EEG) [7], cerebrospinal fluid (CSF) analysis [8], and in-person neuropsychological assessments [9], have been widely investigated. However, these approaches are often resource-intensive, invasive, and costly [10, 11]. Consequently, there is increasing interest in developing non-invasive, cost-effective, and scalable methods for early NCD detection.
@@
One promising approach is the automatic analysis of speech and language, as communicative impairments are often among the earliest symptoms of cognitive decline [12]. These impairments may include temporal changes in spontaneous speech [13], word finding and retrieval difficulties [14], verbal dysfluencies [15], and structural and thematic organization difficulties [16, 17]. Such changes in speech patterns and language use provide valuable insights into underlying cognitive deficits and serve as potential digital biomarkers for early NCD detection. In light of this, various spoken cognitive tests, such as counting [18] and naming [19], have been designed to identify spoken language correlates for NCD detection. Among these, the visual-stimulated narrative (VSN) task [20, 21], such as describing a single image or telling a story based on image sequences, stands out due to its ability to elicit rich and spontaneous speech through visual stimuli within a naturalistic and holistic framework. Specifically, the VSN task involves cognitive-linguistic processes such as visual perception, verbal description, event sequencing, and narrative integration, thereby engaging multiple cognitive domains, including attention, visual comprehension, language fluency, executive planning, memory, and social cognition [22, 23, 24, 25, 26, 27, 28]. Therefore, it holds strong promise for comprehensively assessment of cognitive abilities for NCD detection [29, 12].
@@
Approaches to the analysis of VSNs can be categorized into two distinct yet complementary types [30]. First, the mainstream approach is microstructural analysis, which examines detailed, localized narrative elements, such as linguistic (e.g., lexicon, semantics, syntax) [31, 32] and acoustic (e.g., prosody, phonology) [33, 34] units, primarily reflecting bottom-up (exogenous, stimulus-driven) cognitive processing. In contrast, macrostructural analysis evaluates the overarching organization of the narrative units, such as informativeness (overall information accuracy and completeness), coherence (logical flow of events or concepts) and text-visual consistency (alignment with visual stimuli [^1] This includes both depiction per image and connection across images, bridging informativeness and coherence.[35], generally representing top-down (endogenous, concept-driven) cognitive integration.
@@
Clinical research indicates that individuals with NCDs often exhibit greater impairment in macrostructuring than microstructuring. While microstructural deficits are associated with basic language abilities (perception and production), macrostructural deficits more directly reflect impairments in higher-order cognitive abilities (attention, memory, organization, executive function), manifesting as disrupted topic maintenance, incoherent sequencing, and inconsistency with visual stimuli [30]. Despite this, macrostructural analysis remains underexplored in NCD narratives.
@@
This research gap stems largely from the abstract and conceptual nature of macrostructural features, which makes them inherently difficult to quantify. Traditional approaches to quantify coherence rely on manual assessments, which are subjective and labor-intensive [36]. Recent advances in natural language processing (NLP) have introduced automated methods, such as computing cosine similarities between consecutive textual embeddings or employing large language models (LLMs) as evaluators [37, 38]. However, these methods often fall short for NCD narratives. Embedding-based similarities may erroneously interpret frequent repetitions or fillers as signs of high cohesion, while failing to capture the dynamic evolution of narrative topics over time. LLM-generated scores lack interpretability due to their untraceable decision pathways, undermining their reliability in clinical applications.
@@
To quantify the consistency between narratives and visual stimuli, recent studies employed text-visual pretrained models to measure the highest relevance scores between image regions and sentences [39] or used pattern-matching methods to compute narrative similarity based on fixed references [40, 41]. However, these methods have limitations: local text-image relevance focuses on region-level consistency and fails to capture global narrative coherence or its temporal evolution; meanwhile, pattern-matching approaches depend on predefined references, rendering them inadequate for open-ended, creative tasks such as storytelling. Moreover, these approaches cannot model how high-level narrative structures – such as topics and text-visual alignment – unfold over time in response to sequential visual input. Therefore, there is a need for automated methods to capture the temporal dynamics of narrative content and their alignment with evolving visual stimuli.
@@
This work investigated specific cognitive and linguistic challenges by analyzing topical shifts, their temporal dynamics, and the consistency between narratives and corresponding visual stimuli. We proposed two novel approaches to macrostructural narrative analysis: Dynamic Topic Models (DTM) for temporal analysis of topical evolution; and Text-Image Temporal Alignment Model (TITAN) for cross-modal (i.e., stimulus images and spoken narratives) alignment analysis. These approaches aim to reveal cognitive deficits by identifying narrative impairments associated with NCDs, thereby enhancing our understanding of their impact on communication and cognition. Our analysis was conducted primarily on the Cantonese CU-MARVEL-RABBIT corpus, with the English ADReSS [42] and ADReSSo [43] corpora included for cross-comparison with existing studies. Experimental results on the CU-MARVEL-RABBIT corpus demonstrated the effectiveness of dynamic topic consistency as a macrostructure metric (F1=0.65, AUC=0.78). The proposed TITAN model achieved superior performance (F1=0.72, AUC=0.81) in classification tasks, outperforming models based on handcrafted microstructure and macrostructure feature sets. Moreover, comparative experiments further demonstrated the effectiveness of the proposed dynamic macrostructure modeling approaches for NCD detection (F1=0.8889 for ADReSS and F1=0.8504 for ADReSSo).
@@
The paper is organized as follows: Section II reviews related works; Section III introduces the corpora used in this study; Section IV illustrates the proposed approaches; Section V presents our experimental settings and results; Section VI analyzes and discusses experimental results; and Section VII concludes this paper and outlines directions for future works.
@@ II Related Work
Extensive research has been conducted in the domain of narrative-based NCD detection, leveraging various methodologies to identify biomarkers and assess cognitive impairments. This section reviews pertinent literature encompassing both microstructural and macrostructural narrative analyses, collectively laying the groundwork for more comprehensive narrative assessment in NCD detection. The term “microstructure” refers to fine-grained elements of narrative production, including lexicon, syntax, semantics, phonetics, acoustics, and other smaller linguistic units. In contrast, the term “macrostructure” refers to higher-level aspects of narrative construction, such as global coherence, thematic development, and the overall organization of narratives. Both microstructural and macrostructural analyses are crucial for a comprehensive understanding of cognitive deficits reflected in NCD narratives. However, existing studies are limited in their ability to capture the full spectrum of narrative features, particularly the dynamic aspects of narrative construction. By identifying these limitations, this review highlights the gaps that the proposed methods aim to fill – particularly through the incorporation of temporal dynamics and macrostructural consistency to advance the study of NCD narratives.
@@ II-A Microstructural Analysis in NCD Detection
Individuals with NCDs exhibit speech deficits that affect spoken language across both acoustic and linguistic dimensions. Prior research has explored the various aspects of these impairments in NCD speech, encompassing prosody (e.g., increased dysfluencies and longer, more frequent pauses), phonetics (e.g., abnormal Mel-Frequency Cepstral Coefficients patterns), phonology (e.g., higher rates of word mispronunciations), lexicon (e.g., reduced type-token ratio and elevated stopword ratio), syntax (e.g., reduced parse-tree depth), and semantics (e.g., abnormal patterns in word- and sentence-embeddings) [31, 44, 33, 45, 46, 13, 47, 48, 32, 49, 50, 40, 51, 52, 53, 54]. These microstructural features are essential for understanding the basic language abilities of individuals with NCDs and provide a foundation for narrative analyses in this domain.
Some recent efforts have focused on identifying highly correlated and significant biomarkers of NCDs from the aforementioned features [55, 56, 57, 58, 59]. For example, Cho et al. [59] linked higher CSF p-tau [^2] CSF p-tau: Cerebrospinal fluid phosphorylated tau, a NCD-related biomarker indicating the burden of neurofibrillary tangle pathology. levels (indicating greater severity of NCD pathology) to specific acoustic and linguistic features: lower ratios of prepositions and nouns, higher pause rates, and shorter speech segment durations. Wang et al. [54] proposed several features based on structured clinical rating scales, such as dysfluencies, repetition, content relevance, syntactic structure and irrelevant information. Such microstructural analysis provides valuable insights into local features, but more comprehensive information can be offered through macrostructural approaches that examine broader narrative structure and contexts. Hence, this work extends prior research by examining the macrostructure of VSNs, including topic evolution and cross-modal consistency.
@@ II-B Macrostructural Analysis in NCD Detection
While microstructural analysis highlights specific language production impairments associated with NCDs, macrostructural analysis is equally important as it reveals the integrity of higher-order cognitive processes such as attention, working memory, planning, and organization. This macrostructural perspective is essential for understanding how NCDs affect the construction of coherent and well-organized narratives [60, 61, 62, 30].
Macrostructural analysis in the literature generally focuses on measures of cohesion (the use of linguistic devices for structural unity) and coherence (logical flow and overall consistency). Individuals with NCDs tend to produce narratives characterized by wrong co-referencing, inconsistencies in maintaining the main topic, uncompleted or irrelevant information, and disordered events [63, 64, 65, 66, 67].
Traditional macrostructural features were typically extracted manually based on expert annotations, such as readability [68], relationship between a proposition and previous one or the global topic [69], or connection between events, macro-propositions and main information units [70].
Recent advances leverage machine learning and LLMs for automatic analyze macrostructure of VSNs. Some studies have quantified global semantic coherence through cosine similarity between adjacent utterance embeddings [37, 47], while others have assessed pragmatic coherence via reference-based metrics such as BLEU and METEOR [41]. More recent work has adopted LLM evaluators to automatically predict macro-descriptors [38].
However, as mentioned in Section I, these methods are unsuitable for NCD spoken narratives: similarity scores are confounded by frequent repetitions, reference-based metrics (BLEU, METEOR) are inapplicable due to non-standardized references, and LLM-based evaluation fails to meet interpretability requirements. Additionally, limited research has investigated the temporal dynamics evolution of topics and the cross-modal alignment between visual stimuli and speech in VSN tasks. These gaps underscore the need for more advanced approaches to evaluate NCD narratives from both macrostructural and temporal perspectives, capable of capturing the interplay between evolving topics and the integration of visual and spoken elements.
@@ III Corpora
This study is primarily conducted based on the Cantonese CU-MARVEL-RABBIT (an acronym for Cognitive Assessment Using Machine LeArning EmpoweRed VoicE AnaLysis – Rabbit Story Part) Corpus, with the English ADReSS and ADReSSo Corpora [42, 43] included for comparative purposes. The CU-MARVEL-RABBIT corpus comprises narratives elicited through the Rabbit Storytelling Task – a 15-frame cartoon narrative prompt depicting a girl and her dog searching for a rabbit while encountering various animals and adventures (similar to “Frog, Where Are You?” [71], but with fewer frames and more vibrant visuals). Similarly, the ADReSS and ADReSSo corpora contain speech recordings from the Cookie Theft Picture Description Task, which is based on a single static image depicting two children stealing cookies while a mother is distracted washing dishes [72].
III-A The ADReSS & ADReSSo Corpora
@@
The ADReSS [42] and ADReSSo [43] datasets, developed by the University of Pittsburgh, contain spontaneous speech recordings from elderly English-speaking adults in the Cookie Theft Picture Description task [72]. This task aims to assess participants’ ability to recognize the elements in the picture and produce a coherent narrative based on their spatial relationships and implied events.
The ADReSS corpus consists of audio recordings and associated transcripts with 156 participants, which is partitioned into balanced training (54 HCs and 54 NCDs) and testing (24 HCs and 24 NCDs) subsets. In contrast, the ADReSSo corpus includes a larger number of samples (237 participants), but does not provide any textual transcripts. It is partitioned into a training subset with 79 HCs and 87 NCDs, and a test subset comprising 36 HCs and 35 NCDs.
III-B The CU-MARVEL-RABBIT-RABBIT Corpus